# Efficient Fine-Tuning and Concept Suppression for Pruned Diffusion Models

### v1

@@ Abstract
Recent advances in diffusion generative models have yielded remarkable progress. While the quality of generated content continues to improve, these models have grown considerably in size and complexity. This increasing computational burden poses significant challenges, particularly in resource-constrained deployment scenarios such as mobile devices. The combination of model pruning and knowledge distillation has emerged as a promising solution to reduce computational demands while preserving generation quality. However, this technique inadvertently propagates undesirable behaviors, including the generation of copyrighted content and unsafe concepts, even when such instances are absent from the fine-tuning dataset. In this paper, we propose a novel bilevel optimization framework for pruned diffusion models that consolidates the fine-tuning and unlearning processes into a unified phase. Our approach maintains the principal advantages of distillation—namely, efficient convergence and style transfer capabilities—while selectively suppressing the generation of unwanted content. This plug-in framework is compatible with various pruning and concept unlearning methods, facilitating efficient, safe deployment of diffusion models in controlled environments.
@@ 1 Introduction
While distillation enhances convergence speed and preserves the expressive power of the original model, it introduces inherent risks. Diffusion models are known to generate copyrighted or inappropriate (e.g., NSFW) content [36, 44], prompting ongoing efforts to filter out such content [8, 20, 41, 55, 50, 33] without the significant expense of retraining on a filtered dataset. It might seem intuitive that removing unsafe concepts from the fine-tuning dataset would naturally filter out such content. However, this assumption does not necessarily hold. Even when such content is excluded from the fine-tuning dataset, unsafe concepts still persist in a pruned diffusion model, particularly when distillation is used during the fine-tuning process (See Figs. 1 and 2). This poses a critical challenge for deploying diffusion models in controlled environments, where generating these types of outputs is unacceptable.
@@ 2.1 Efficient Diffusion Architectures
Numerous studies have targeted efficient architecture design for diffusion models [57, 51, 25]. For instance, MobileDiffusion [57] applies empirical adjustments based on performance metrics from MS-COCO [26], while SnapFusion [25] focuses on finding optimal architectures specifically for text-to-image (T2I) models. Pruning-based approaches attempt to eliminate non-essential layers or blocks of a pre-trained diffusion model. SPDM [5] evaluates the importance of weights via Taylor expansion, removing lower-value weights, while BK-SDM [19] streamlines the U-Net by eliminating blocks that minimally impact generation quality. APTP [10] employs a dynamic, prompt-based pruning strategy for T2I models, adjusting resources according to prompt complexity in a Mixture of Experts setting. Some pruning-based approaches utilize distillation techniques—both knowledge distillation [15] and feature distillation [38]—to retrain the pruned model [19, 10, 12, 22, 53], restoring its generative abilities.
@@ 2.2 Concept Editing and Unlearning in Diffusion Models
Diffusion model unlearning (MU) methods [49, 56, 30, 4, 13], by contrast, rely on a “forgetting” dataset to remove specific information. These methods employ optimization approaches like dual problem formulations [49], generative replay for reinforcing retention [13], and saliency-based fine-tuning masks [4]. Unlike these approaches, ConceptPrune [2] offers a training-free pruning method to remove regions responsible for undesired outputs.
@@ 3.1 Diffusion Models
$\mathcal{L}_{D}^{\text{Diff}}(\theta)=\mathbb{E}_{x_{0},t,c,\epsilon}[w(t)\|\epsilon_{\theta}(x_{t},t,c)-\epsilon\|^{2}],$ (1)
@@ 3.2 Model Pruning & Distillation
$\min_{\theta_{\text{pruned}}}|L(\theta_{\text{pruned}})-L(\theta)|,\quad\text{s.t.}\quad\|\theta_{\text{pruned}}\|_{0}\leq R,$ (2)
@@
After pruning, the model usually loses its high quality generation capabilities and requires an additional fine-tuning stage.
Previous approaches to retraining pruned diffusion models [19, 10, 53], sometimes leverage additional objectives such as output distillation and feature distillation [15, 38] together with the original denoising objective (Eq. 1). These objectives encourage the pruned model (student) to match the behavior of the original model (teacher) across both output predictions and intermediate feature representations. Specifically, the distillation objectives are defined as:
@@ 3.3 Concept Unlearning
Following prior work [8, 9, 20], we define concept unlearning for a pre-trained generative model $p_{\theta}(x,c)$ as the task of preventing the generation of a specific concept $c$. A natural question then arises: what should be generated in place of this omitted concept? Previous approaches often guide the model to generate samples conditioned on a related anchor concept, denoted by $c^{\prime}$. The anchor concept $c^{\prime}$ could be a similar concept to the target concept $c$—for example, Anchor: ”Cat” vs. Target: ”Grumpy Cat” [20]. Alternatively, $c^{\prime}$ could represent a ”null” concept, which encourages the model to revert to generating samples that resemble the unconditional outputs of the pre-trained model [18]. The unlearning objective would then be
@@ 4 Method
$\min_{\theta_{\text{pruned}}}\mathcal{L^{\text{ft}}}:=\mathcal{L}^{\text{Diff}}+\lambda^{\text{OutKD}}\mathcal{L}^{\text{OutKD}}+\lambda^{\text{FeatKD}}\mathcal{L}^{\text{FeatKD}},$ (8)
@@
Figure 3: Why can a two-stage approach (fine-tuning followed by forgetting) be suboptimal? If fine-tuning yields $\hat{\theta}$, initializing the concept unlearning parameters with $\hat{\theta}$ and optimizing the concept unlearning loss (Eq. 5) results in $\theta^{\prime}$, which is suboptimal for both fine-tuning the pruned model and for concept unlearning. In contrast, our bilevel method, defined in Eq. 9, produces the optimal solution $\theta^{*}$, achieving better performance for both fine-tuning and unlearning.
@@
$\begin{split}\min_{\theta_{\text{pruned}}}&\mathbb{E}_{x_{0},\epsilon,t,c,c^{\prime}}\|\epsilon_{\theta}(x_{t},t,c^{\prime})-\epsilon_{\theta_{\text{pruned}}}(x_{t},t,c)\|^{2},\\ s.t.\ &\mathcal{L^{\text{ft}}}(\theta_{\text{pruned}})-\inf_{\vartheta}\mathcal{L^{\text{ft}}}(\vartheta)\leq 0.\end{split}$ (10)
@@
$\min_{\theta_{\text{pruned}}}\mathcal{L^{\text{penalized}}}(\theta_{\text{pruned}}),$ (11)
@@ 5.1 Effect of Distillation and Pruning
Figure 4: Effect of Distillation: Adding output- and feature-level distillation to the fine-tuning process of a pruned model significantly accelerates convergence, especially in resource-constrained settings.
@@
Figure 5: Effect of Pruning: Initializing the weights of a pruned model with those of the teacher model enables significantly faster convergence compared to random initialization, making pruning combined with distillation an excellent choice for training a small diffusion model.
@@ 5.2.2 Style Removal Results
CLIP [35] ($\downarrow$) CP [2] ($\uparrow$) CSD [45] ($\downarrow$) FID ($\downarrow$) CLIP ($\uparrow$)
@@
Following the evaluations from [2], we assess the effectiveness of our approach and the baselines in removing the styles of three artists—Vincent Van Gogh, Pablo Picasso, and Claude Monet—whose styles are effectively replicated by Stable Diffusion [8]. Using a dataset of 50 prompts for each artist’s style, we report the CLIP similarity [35] between generated samples and the prompts, as well as a stricter CLIP score that penalizes the unlearned model when its generated samples are more similar to the prompts than those of the original model from Chavhan et al. [2]. We also report the CSD Score [45], a metric specifically proposed for measuring style similarity. The standard CLIP-based metrics quantify general alignment between the generated samples and the prompts. In contrast, the CSD score specifically targets style similarity, allowing for the assessment of stylistic nuances that may be missed by CLIP-based metrics. These metrics together provide a more comprehensive evaluation by capturing different aspects of the unlearning process.
@@
In  Fig. 6, we present qualitative results illustrating the concept erasure capabilities of our proposed method compared to the baselines. Our bilevel method effectively removes the desired concept entirely. Note the subtle similarities between the generated images of the baselines and the original model in the first three rows, where some leakage indicates that the model still retains aspects of the removed style. For instance, the prompt used for the samples in the style of Van Gogh is “The Weaver by Vincent van Gogh”. Although the original Van Gogh painting depicts a woman, the prompt does not specify the subject’s gender. However, baseline methods continue to generate an image of a woman, suggesting that residual elements of Van Gogh’s painting persist in them even after removal. In contrast, our method generates high-quality images without any trace of the erased styles. Furthermore, our method excels at preserving other non-removed styles, as evidenced by the comparison between images from the original model and our method in the last three rows. The baselines, however, exhibit style interference, altering other non-removed styles as well.
@@ 5.2.3 Explicit Content Removal Results
MMA ($\uparrow$) Ring-a-Bell ($\uparrow$)
Distilled Model + ESD [8] 93.70 77.27
Distilled Model + UCE [9] 88.57 76.14
Distilled Model + CP [2] 94.12 97.72
Ours 91.60 94.32
Table 2: Comparison of our bilevel method with baseline removal methods on adversarial NSFW prompts: The values indicate the resilience of each method to adversarial prompts. While our method does not outperform all baselines, it demonstrates solid performance on these challenging prompts.
When interpreting the results in Fig. 7 and Tab. 2, it is important to consider generation quality on other unaffected concepts, as shown in Tab. 3. Our method performs comparably to the baselines on NSFW removal, with significantly less impact on generation quality judged by FID and CLIP scores on 5000 validation prompts of MS-COCO. Our method offers a favorable trade-off between explicit content reduction and output quality.
COCO
FID ($\downarrow$) CLIP ($\uparrow$)
Distilled Model + ESD [8] 32.47 28.57
Distilled Model + UCE [9] 41.55 26.60
Distilled Model + CP [2] 29.56 29.45
Ours 26.80 29.94
Table 3: Quantitative results demonstrating the model’s ability to retain styles and concepts that were not targeted for removal by the NSFW removal method. Our bilevel method does not impact the generation capabilities of the model.
@@ 6 Conclusion
In this paper, we introduced a bilevel optimization framework for fine-tuning pruned diffusion models, addressing the dual challenge of restoring generative quality while effectively suppressing unwanted concepts. Our approach leverages a synergistic optimization structure, where the lower-level optimization performs standard fine-tuning with denoising and distillation loss minimization to maintain the expressive power of the pruned model, and the upper-level optimization selectively unlearns specific concepts. Through comprehensive experiments, we demonstrated that our framework outperforms baseline methods in both generative quality and concept removal tasks, including artist style erasure and NSFW content suppression.
Our proposed method is versatile, acting as a plug-in compatible with various pruning and concept unlearning techniques, making it suitable for adapting pruned diffusion models to diverse controlled environments. This flexibility is especially valuable for deploying generative models in settings where ethical and practical constraints demand careful control over the content.
@@ C.1.3 Concept Removal Experimental Settings
  • ESD [8]: ESD is a method for erasing concepts from text-to-image diffusion models by fine-tuning the model weights using negative guidance. The goal is to reduce the probability of generating images associated with a specific concept, represented by $P_{\theta}(x)\propto\frac{P_{\theta^{*}}(x)}{P_{\theta^{*}}(c|x)^{\eta}}$, where $P_{\theta}(x)$ is the distribution of the edited model, $P_{\theta^{*}}(x)$ is the distribution of the original model, $c$ is the concept to be erased, and $\eta$ is a scaling factor. By manipulating the gradient of the log probability, the authors arrive at a modified score function: $\epsilon_{\theta}(x_{t},c,t)\leftarrow\epsilon_{\theta^{*}}(x_{t},t)-\eta[\epsilon_{\theta^{*}}(x_{t},c,t)-\epsilon_{\theta^{*}}(x_{t},t)]$. This function guides the model away from the undesired concept during fine-tuning. The method uses the model’s existing knowledge of the concept to generate training samples, eliminating the need for additional data. ESD offers two variations: ESD-x for prompt-specific erasure, such as artistic styles, and ESD for global erasure, such as nudity. Similar to the original paper we remove ”Van Gogh”, ”Claude Monet”, and ”Picasso” from the diffusion model for artist erasure, and remove ”nudity” for explicit content erasure. This process uses the AdamW [31] optimizer with a learning rate of $0.00001$, and a negative guidance $\eta=1$. The model is trained for 1000 iterations to remove the concept. For artist style and explicit content removal we pick ”ESD-x” and ”ESD-u”, respectively.
@@
  • UCE [9]: UCE is a method for editing multiple concepts in text-to-image diffusion models without retraining. UCE works by directly modifying the attention weights of the model in a closed-form solution, making it efficient and scalable. The method aims to address various safety issues such as bias, copyright infringement, and offensive content, which previous methods have tackled separately. UCE modifies the cross-attention weights, denoted as $W$, to minimize the difference between the model’s output for the concepts to edit, $c_{i}$, and their desired target output, $v^{*}_{i}$. This is achieved by minimizing the objective function: $\sum_{c_{i}inE}||Wc_{i}-v^{*}_{i}||^{2}_{2}+\sum_{c_{j}\in P}||Wc_{j}-W^{old}c_{j}||^{2}_{2}$ where E represents the set of concepts to edit and P represents the set of concepts to preserve. This formula ensures that the model’s output for the edited concepts is steered towards the desired target, while preserving the output for the concepts that should remain unchanged. Identical to the original setting of the paper, we remove ”Van Gogh”, ”Claude Monet”, and ”Pablo Picasso” for artist erasure and guide them towards ”art”. We remove ”nudity” for explicit content removal and guide them towards ”person”. Other hyperparameters are identical to the values set in their training code.
@@
  • ConceptPrune [2] ConceptPrune is a method for removing unwanted concepts from pre-trained text-to-image diffusion models without any retraining. This is achieved by pruning or zeroing out specific neurons within the model’s feed-forward networks that are identified as being responsible for generating the unwanted concept. This method is inspired by the observation that certain neurons in neural networks specialize in specific concepts. ConceptPrune first Identifies skilled neurons by analyzing the activation patterns of neurons in response to prompts with and without the unwanted concept and then prunes them. For ConceptPrune, we set skill ratio to $0.01$. We remove ”Van Gogh”, ”Claude Monet”, and ”Picasso” from the diffusion model for artist erasure, and remove ”nudity” for explicit content erasure. Other hyperparameters are identical to best settings in their released code.
@@
C.2 Figure Prompts

### v2

@@ Abstract
Recent advances in diffusion generative models have yielded remarkable progress. While the quality of generated content continues to improve, these models have grown considerably in size and complexity. This increasing computational burden poses significant challenges, particularly in resource-constrained deployment scenarios such as mobile devices. The combination of model pruning and knowledge distillation has emerged as a promising solution to reduce computational demands while preserving generation quality. However, this technique inadvertently propagates undesirable behaviors, including the generation of copyrighted content and unsafe concepts, even when such instances are absent from the fine-tuning dataset. In this paper, we propose a novel bilevel optimization framework for pruned diffusion models that consolidates the fine-tuning and unlearning processes into a unified phase. Our approach maintains the principal advantages of distillation—namely, efficient convergence and style transfer capabilities—while selectively suppressing the generation of unwanted content. This plug-in framework is compatible with various pruning and concept unlearning methods, facilitating efficient, safe deployment of diffusion models in controlled environments. Code is available here.
@@ 1 Introduction
While distillation enhances convergence speed and preserves the expressive power of the original model, it introduces inherent risks. Diffusion models are known to generate copyrighted or inappropriate (e.g., NSFW) content [39, 47], prompting ongoing efforts to filter out such content [9, 23, 44, 59, 53, 36, 58, 20] without the significant expense of retraining on a filtered dataset. It might seem intuitive that removing unsafe concepts from the fine-tuning dataset would naturally filter out such content. However, this assumption does not necessarily hold. Even when such content is excluded from the fine-tuning dataset, unsafe concepts still persist in a pruned diffusion model, particularly when distillation is used during the fine-tuning process (See Figs. 1 and 2). This poses a critical challenge for deploying diffusion models in controlled environments, where generating these types of outputs is unacceptable.
@@ 2.1 Efficient Diffusion Architectures
Numerous studies have targeted efficient architecture design for diffusion models [61, 54, 28]. For instance, MobileDiffusion [61] applies empirical adjustments based on performance metrics from MS-COCO [29], while SnapFusion [28] focuses on finding optimal architectures specifically for text-to-image (T2I) models. Pruning-based approaches [11] attempt to eliminate non-essential layers or blocks of a pre-trained diffusion model. SPDM [6] evaluates the importance of weights via Taylor expansion, removing lower-value weights, while BK-SDM [22] streamlines the U-Net by eliminating blocks that minimally impact generation quality. APTP [12] employs a dynamic, prompt-based pruning strategy for T2I models, adjusting resources according to prompt complexity in a Mixture of Experts setting. Some pruning-based approaches utilize distillation techniques—both knowledge distillation [17] and feature distillation [41]—to retrain the pruned model [22, 12, 14, 25, 56], restoring its generative abilities.
@@ 2.2 Concept Editing and Unlearning in Diffusion Models
Diffusion model unlearning (MU) methods [52, 60, 33, 5, 15, 58], by contrast, rely on a “forgetting” dataset to remove specific information. These methods employ optimization approaches like dual problem formulations [52], generative replay for reinforcing retention [15], and saliency-based fine-tuning masks [5]. Unlike these approaches, ConceptPrune [2] offers a training-free pruning method to remove regions responsible for undesired outputs.
@@ 3.1 Diffusion Models
$\mathcal{L}_{D}^{\text{Diff}}(\theta)=\mathbb{E}_{x_{0},t,c,\epsilon}[w(t)\|\epsilon_{\theta}(x_{t},t,c)-\epsilon\|^{2}],\vspace{-5pt}$ (1)
@@ 3.2 Model Pruning & Distillation
$\vspace{-5pt}\min_{\theta_{\text{pruned}}}|L(\theta_{\text{pruned}})-L(\theta)|,\quad\text{s.t.}\quad\|\theta_{\text{pruned}}\|_{0}\leq R,$ (2)
@@
After pruning, the model usually loses its high quality generation capabilities and requires an additional fine-tuning stage. Previous approaches to retraining pruned diffusion models [22, 12, 56], sometimes leverage additional objectives such as output distillation and feature distillation [17, 41] together with the original denoising objective (Eq. 1). These objectives encourage the pruned model (student) to match the behavior of the original model (teacher) across both output predictions and intermediate feature representations. Specifically, the distillation objectives are defined as:
@@ 3.3 Concept Unlearning
Following prior work [9, 10, 23], we define concept unlearning for a pre-trained generative model $p_{\theta}(x,c)$ as the task of preventing the generation of a specific concept $c$. A natural question then arises: what should be generated in place of this omitted concept? Previous approaches often guide the model to generate samples conditioned on a related anchor concept, denoted by $c^{\prime}$. The anchor concept $c^{\prime}$ could be a similar concept to the target concept $c$—for example, Anchor: "Cat" vs. Target: "Grumpy Cat" [23]. Alternatively, $c^{\prime}$ could represent a "null" concept, which encourages the model to revert to generating samples that resemble the unconditional outputs of the pre-trained model [21]. The unlearning objective would then be
@@ 4 Method
$\min_{\theta_{\text{pruned}}}\mathcal{L^{\text{ft}}}:=\mathcal{L}^{\text{Diff}}+\lambda^{\text{OutKD}}\mathcal{L}^{\text{OutKD}}+\lambda^{\text{FeatKD}}\mathcal{L}^{\text{FeatKD}},\vspace{-5pt}$ (8)
@@
Figure 3: Why can a two-stage approach (fine-tuning followed by forgetting) be suboptimal? If fine-tuning yields $\hat{\theta}$, initializing the concept unlearning parameters with $\hat{\theta}$ and optimizing the concept unlearning loss (Eq. 5) results in $\theta^{\prime}$, which is suboptimal for both fine-tuning and for concept unlearning. In contrast, our bilevel method (Eq. 9) produces the optimal solution $\theta^{*}$, achieving better performance for both fine-tuning and unlearning.
@@
$\begin{split}\min_{\theta_{\text{pruned}}}&\mathbb{E}_{x_{0},\epsilon,t,c,c^{\prime}}\|\epsilon_{\theta}(x_{t},t,c^{\prime})-\epsilon_{\theta_{\text{pruned}}}(x_{t},t,c)\|^{2},\\ s.t.\ &\mathcal{L^{\text{ft}}}(\theta_{\text{pruned}})-\inf_{\vartheta}\mathcal{L^{\text{ft}}}(\vartheta)\leq 0.\end{split}\vspace{-5pt}$ (10)
@@
$\min_{\theta_{\text{pruned}}}\mathcal{L^{\text{penalized}}}(\theta_{\text{pruned}}),\vspace{-5pt}$ (11)
@@ 5.1 Effect of Distillation and Pruning
Figure 4: Effect of Distillation: Adding distillation to the fine-tuning process of a pruned model significantly accelerates convergence, especially in resource-constrained settings.
@@
Figure 5: Effect of Pruning: Pruning enables significantly faster convergence compared to random initialization, making it an excellent choice for training a small diffusion model.
@@ 5.2.2 Style Removal Results
CLIP ($\downarrow$) CP ($\uparrow$) CSD ($\downarrow$) FID ($\downarrow$) CLIP ($\uparrow$)
@@
Following the evaluations from [2], we assess the effectiveness of our approach and the baselines in removing the styles of three artists—Vincent Van Gogh, Pablo Picasso, and Claude Monet—whose styles are effectively replicated by Stable Diffusion [9]. Using a dataset of 50 prompts for each artist’s style, we report the CLIP similarity [38] between generated samples and the prompts, as well as a stricter CLIP score that penalizes the unlearned model when its generated samples are more similar to the prompts than those of the original unpruned model (introduced in  [2]). We also report the CSD Score [48], a metric specifically proposed for measuring style similarity. The standard CLIP-based metrics quantify general alignment between the generated samples and the prompts. In contrast, the CSD score specifically targets style similarity, allowing for the assessment of stylistic nuances that may be missed by CLIP-based metrics. These metrics together provide a more comprehensive evaluation by capturing different aspects of the unlearning process.
@@
In  Fig. 6, we present qualitative results comparing the concept erasure capabilities of our proposed method and the baselines. Our bilevel method effectively removes the desired concept entirely. Note the subtle similarities between the generated images of the baselines and the original model in the first three rows, where some leakage indicates that the model still retains aspects of the removed style. For instance, the prompt used for the samples in the style of Van Gogh is “The Weaver by Vincent van Gogh". Although the original Van Gogh painting depicts a woman, the prompt does not specify the subject’s gender. However, baseline methods continue to generate an image of a woman, suggesting that residual elements of Van Gogh’s painting persist in them even after removal. In contrast, our method generates high-quality images without any trace of the erased styles. Moreover, our method excels at preserving other non-removed styles, as evidenced by the comparison between images from the original model and our method in the last three rows. The baselines, however, exhibit style interference, altering other non-removed styles as well.
@@ 5.2.3 Explicit Content Removal Results
Method NSFW Removal Bench. COCO
MMA ($\uparrow$) Ring-a-Bell ($\uparrow$) FID ($\downarrow$) CLIP ($\uparrow$)
Distilled Model + ESD [9] 93.70 77.27 32.47 28.57
Distilled Model + UCE [10] 88.57 76.14 41.55 26.60
Distilled Model + CP [2] 94.12 97.72 29.56 29.45
Ours (ESD) 91.60 94.32 26.80 29.94
Table 2: Comparison of our method with two-stage baselines. Our method shows competitive NSFW removal while achieving superior generation quality on concepts not targeted for removal.
When interpreting the results in Fig. 7 and Tab. 2, it is important to consider generation quality on other unaffected concepts. As shown in Tab. 2. Our method performs comparably to the baselines on NSFW removal, with significantly less impact on generation quality judged by FID and CLIP scores on 5000 validation prompts of MS-COCO. Our method offers a favorable trade-off between explicit content reduction and output quality.
As mentioned before, our framework is compatible with any unlearning technique. Our baseline is a two stage distillation + unlearning framework rather than a specific unlearning technique. The objective of our experiments was to demonstrate how our bilevel method even when paired with a basic unlearning approach like ESD [9] can outperform two stage baselines with powerful unlearning methods. Paired with a more powerful unlearning method, our approach can outperform the baselines (See Sec. C.2).
5.3 Continual Unlearning
Our method produces an efficient, and unlearned checkpoint that retains its original generation capability. Any unlearning approach can be applied on the result if unlearning of another concept is needed. Nonetheless, the reasoning outlined in Sec. 4 also applies to continual unlearning—i.e., a two-stage process of first unlearning $c_{1}$ and then unlearning $c_{2}$ may be suboptimal. However, we believe the degree of suboptimality would be less severe than the challenge addressed in this paper—i.e. restoring the generative capability of a pruned model while suppressing unwanted concepts. This presents a compelling direction for future research.
@@ 6 Conclusion
In this paper, we introduce a bilevel optimization framework for fine-tuning pruned diffusion models, addressing the dual challenge of restoring generative quality while effectively suppressing unwanted concepts. Our experiments show that our framework outperforms two-stage baseline methods in both generative quality and concept removal tasks, including artist style erasure and NSFW content suppression. Our proposed method is compatible with various pruning and concept unlearning techniques. This method is especially valuable for deploying generative models in settings where ethical and practical constraints demand careful control over the content.
@@ C.1.3 Concept Removal Experimental Settings
  • ESD [9]: ESD is a method for erasing concepts from text-to-image diffusion models by fine-tuning the model weights using negative guidance. The goal is to reduce the probability of generating images associated with a specific concept, represented by $P_{\theta}(x)\propto\frac{P_{\theta^{*}}(x)}{P_{\theta^{*}}(c|x)^{\eta}}$, where $P_{\theta}(x)$ is the distribution of the edited model, $P_{\theta^{*}}(x)$ is the distribution of the original model, $c$ is the concept to be erased, and $\eta$ is a scaling factor. By manipulating the gradient of the log probability, the authors arrive at a modified score function: $\epsilon_{\theta}(x_{t},c,t)\leftarrow\epsilon_{\theta^{*}}(x_{t},t)-\eta[\epsilon_{\theta^{*}}(x_{t},c,t)-\epsilon_{\theta^{*}}(x_{t},t)]$. This function guides the model away from the undesired concept during fine-tuning. The method uses the model’s existing knowledge of the concept to generate training samples, eliminating the need for additional data. ESD offers two variations: ESD-x for prompt-specific erasure, such as artistic styles, and ESD for global erasure, such as nudity. Similar to the original paper we remove "Van Gogh", "Claude Monet", and "Picasso" from the diffusion model for artist erasure, and remove "nudity" for explicit content erasure. This process uses the AdamW [34] optimizer with a learning rate of $0.00001$, and a negative guidance $\eta=1$. The model is trained for 1000 iterations to remove the concept. For artist style and explicit content removal we pick "ESD-x" and "ESD-u", respectively.
@@
  • UCE [10]: UCE is a method for editing multiple concepts in text-to-image diffusion models without retraining. UCE works by directly modifying the attention weights of the model in a closed-form solution, making it efficient and scalable. The method aims to address various safety issues such as bias, copyright infringement, and offensive content, which previous methods have tackled separately. UCE modifies the cross-attention weights, denoted as $W$, to minimize the difference between the model’s output for the concepts to edit, $c_{i}$, and their desired target output, $v^{*}_{i}$. This is achieved by minimizing the objective function: $\sum_{c_{i}inE}||Wc_{i}-v^{*}_{i}||^{2}_{2}+\sum_{c_{j}\in P}||Wc_{j}-W^{old}c_{j}||^{2}_{2}$ where E represents the set of concepts to edit and P represents the set of concepts to preserve. This formula ensures that the model’s output for the edited concepts is steered towards the desired target, while preserving the output for the concepts that should remain unchanged. Identical to the original setting of the paper, we remove "Van Gogh", "Claude Monet", and "Pablo Picasso" for artist erasure and guide them towards "art". We remove "nudity" for explicit content removal and guide them towards "person". Other hyperparameters are identical to the values set in their training code.
@@
  • ConceptPrune [2] ConceptPrune is a method for removing unwanted concepts from pre-trained text-to-image diffusion models without any retraining. This is achieved by pruning or zeroing out specific neurons within the model’s feed-forward networks that are identified as being responsible for generating the unwanted concept. This method is inspired by the observation that certain neurons in neural networks specialize in specific concepts. ConceptPrune first Identifies skilled neurons by analyzing the activation patterns of neurons in response to prompts with and without the unwanted concept and then prunes them. For ConceptPrune, we set skill ratio to $0.01$. We remove "Van Gogh", "Claude Monet", and "Picasso" from the diffusion model for artist erasure, and remove "nudity" for explicit content erasure. Other hyperparameters are identical to best settings in their released code.
@@
C.2 More Results
Our framework is compatible with any unlearning technique. Our baseline is a two stage distillation + unlearning framework rather than a specific unlearning technique. The objective of our experiments was to demonstrate how our bilevel method even when paired with a basic unlearning approach like ESD can outperform two stage baselines with powerful unlearning methods.
C.2.1 Unlearn - Prune - Distill Baseline
A natural question would be to compare our method with an alternative baseline where the concepts are first unlearnt from base model and then distillation is done. This baseline requires more resources than our method since unlearning occurs on the unpruned model. However, our analysis applies to this approach too. We conduct an experiment for this baseline (See Tab. 3 rows 1 and 2), and the generation quality is worse. This is expected—applying unlearning to the base model reduces its quality. Using this degraded model for distillation further impacts the performance of the already weaker pruned model.
Method ASR ($\downarrow$) FID ($\downarrow$)
Distilled Model + ESD [9] 62.7 32.47
ESD [9] + Distilled Model 59.8 39.11
Distilled Model + UCE [10] 67.6 41.55
Distilled Model + CP [2] 54.9 29.56
Distilled Model + AdvUnlearn [58] 36.6 36.17
Ours (ESD [9]) 57.0 26.80
Ours (AdvUnlearn [58]) 32.4 26.91
Table 3: Attack Success Rate of adversarial NSFW prompts from  [4] and generation quality on COCO-Val-2017
To show this compatibility with other unlearning methods, we ran two additional experiments: (1) Two-stage pipeline with the recently proposed AdvUnlearn [58] and (2) our bilevel framework with AdvUnlearn [58]. Results in Tab. 3 confirm our method integrates well with AdvUnlearn [58] and achieves superior performance in terms of ASR and FID compared to all baselines.
We provided more samples from our method in Fig. 8.
Figure 8: More visual samples of our bilevel method(with ESD)
C.3 Figure Prompts